OpenStudy (anonymous):

The table weighs 25 kilograms. Find the equivalent weight in pounds. Round off your decimal answer to 2 decimal places. Enter only significant zeros.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

1lb=0.4535kg, ?lb=25kg 25kg=25/4535=55.11

OpenStudy (anonymous):

sorry 25kg=25/0.4535=55.11

OpenStudy (anonymous):

is 55.11 rounded off to the 2nd decimal right.

OpenStudy (anonymous):

@SalvadorV

OpenStudy (mathmale):

Salvador's conversion factor is more accurate than mine, but I trust mine a bit further: 1 kg corresponds to 2.2 lb. Thus, the weight of a 25 Kg table, in lb., would be \[\frac{ 25 kg }{ 1 }*\frac{ 2.2lb }{ 1 kg }\]

OpenStudy (anonymous):

yeah, its to the second decimal

OpenStudy (mathmale):

Try multiplying that out.

OpenStudy (anonymous):

55bl

OpenStudy (anonymous):

lb*

OpenStudy (mathmale):

Notice how close that is to Salvador's result? My conversion factor is an approx., whereas his is much more accurate. Notice how he did round off his answer to 2 decimal places? What will you choose to be your final answer to this problem?

OpenStudy (anonymous):

55.11 than right ?

OpenStudy (mathmale):

Provided that Salvador is right (and I have no reason to believe that he isn't), YES! Be sure you know where and how to obtain the appropriate conversion factor(s); personally, I'd do an internet search for exactly that, or for "kg to pounds conversion."

OpenStudy (anonymous):

thank you mathmale your too smart thanks

OpenStudy (mathmale):

Many thanks! Enjoyed working with you! Bye for now.